OpenStudy (hannahc234):

Check my answer A community group is planning the expansion of a square flower garden in a city park. If each side of the original garden is increased by 3 meters, the new total area of the garden will be 225 square meters. Find the length of each side of the original garden. 15 m 3 m 12 m <---- √12 m

OpenStudy (photon336):

Your answer is correct here is how we can figure this out  \[Area = s^{2} or~(x+3)^{2} \] \[x^{2}+6x+9 = 225 \] \[x^{2}+6x-216 = 0\] \[\frac{ \sqrt{6^{2}-4(1)(-216)} }{ 2 } = \frac{ (-6+ 30) }{ 2 } = 12 m \] \[(12)^{2} =144, (15)^{2} = 225 \]